# Collective steering: Tracer-informed dynamics

### v1

@@ 4 Internal Dynamics from Tracer Trajectories
This section deals with a counterpart of Problem 1. In it, the data of the problem we consider here is in the form of the distributional constraints (4) with $\mathcal{T}_{1}=\{0,1\}$ and the full trajectories of the tracers (5) with $\mathcal{T}_{2}=[0,1]$. Our task is once again to identify the dynamics in (3). If (7) is taken as the cost of transport, then a candidate problem formulation can be stated as follows.
Problem 4.1.
Given $\Sigma_{t\in\mathcal{T}_{1}}\in{\rm Sym}^{+}(n)$ with $\mathcal{T}_{1}=\{0,1\}$ and $Y_{t\in\mathcal{T}_{2}}\in\mathbb{R}^{n\times m}$ with $\mathcal{T}_{2}=[0,1]$, find the minimizer of (7), subject to the constraints (3), (4), and (5).
@@
However, this candidate formulation may not admit a solution, in general; there are cases where $\Phi_{t}$ becomes singular at some point in time, see [8, Example 1]. In order to avert this issue, we modify the cost in Problem 4.1 by augmenting (7) with an added “weighted attention” term (9) that penalizes the $L_{2}$-norm of the gain $K_{t}$. Such an augmentation traces back to R.W. Brockett [5], according to whom, “… the easiest control law to implement is a constant input. Anything else requires some attention. The more frequently the control changes, the more effort it takes to implement it. […], it can be argued that the cost of implementation is linked to the rate at which the control changes with changing values of $x$ [state] and $t$ [time].”
@@
For convenience, we restrict ourselves to a spatial attention term as in (8) or (9). Thereby, we state the adaptation of Problem 4.1 as follows.
Problem 4.2.
Given $\Sigma_{t\in\mathcal{T}_{1}}\in{\rm Sym}^{+}(n)$ with $\mathcal{T}_{1}=\{0,1\}$ and $Y_{t\in\mathcal{T}_{2}}\in\mathbb{R}^{n\times m}$ with $\mathcal{T}_{2}=[0,1]$, find the minimizer of
[…]
@@
A curve $(K_{t},\Phi_{t})_{t\in[0,1]}$ is said to be admissible for Problem 4.2 if $K_{t\in[0,1]}$ is Lebesgue-integrable and $\Phi_{t\in[0,1]}$ is absolutely continuous, and are such that (3) is satisfied (whenever $\dot{\Phi}_{t}$ exists) as well as constraints (4) and (5), and the value of the functional (14) is finite. The following proposition provides a more explicit characterization of the class of admissible curves.
@@
Then, a curve $(K_{t},\Phi_{t})_{t\in[0,1]}$ is an admissible curve for Problem 4.2 if and only if the following four conditions hold. First, the constraint (4) is satisfied, second,
@@
third, there exists an almost everywhere unique Lebesgue integrable curve $R_{t\in[0,1]}\in\mathbb{R}^{n\times(n-m)}$ such that
@@
Proof 4.3.
Let $(K_{t},\Phi_{t})_{t\in[0,1]}$ be an admissible curve for Problem 4.2. Since $\mathcal{T}_{2}=[0,1]$ and $Y_{t\in\mathcal{T}_{2}}$ is continuously differentiable by assumption, direct differentiation of (5) followed by substitution of (3) shows that $K_{t\in[0,1]}$ must satisfy
@@
where $Z_{t\in[0,1]}\in\mathbb{R}^{n\times n}$ is any Lebesgue-integrable curve satisfying $Z_{t}Y_{t}=0$, i.e., the column space of $Y_{t}$ is a subspace of the null space of $Z_{t}$.
To proceed, let $N_{0}\in\mathbb{R}^{(n-m)\times n}$ be any full row-rank matrix such that $N_{0}Y_{0}=0$ and define the curve $N_{t\in[0,1]}$ as the solution to
@@
From the above, it follows that any Lebesgue-integrable curve $Z_{t\in[0,1]}\in\mathbb{R}^{n\times n}$ satisfying the constraint $Z_{t}Y_{t}=0$ must be of the form $Z_{t}=R_{t}N_{t}$, for an almost everywhere unique Lebesgue-integrable curve $R_{t\in[0,1]}\in\mathbb{R}^{n\times(n-m)}$. Finally, since $1\in\mathcal{T}_{2}=[0,1]$, the curve $\Phi_{t\in[0,1]}$ must be such that $\Phi_{1}$ satisfies (16). The converse implication of the proposition follows by direct integration.
Proposition 4 suggests a reformulation of Problem 4.2, replacing condition (5) on the state $\Phi_{t}$ with an explicit constraint on the gain $K_{t}$. The advantage of the new formulation lies in the fact that the search for the optimal curve reduces to searching over a lower dimensional control space, namely, the curve $R_{t\in[0,1]}$.
Problem 4.2
Given $\Sigma_{t\in\mathcal{T}_{1}}\in{\rm Sym}^{+}(n)$ with $\mathcal{T}_{1}=\{0,1\}$ and $Y_{t\in\mathcal{T}_{2}}\in\mathbb{R}^{n\times m}$ with $\mathcal{T}_{2}=[0,1]$, find the minimizer of (14) for $\varepsilon>0$ and subject to the constraints (3), (4), (16), and (17).
Before we proceed with deriving the necessary conditions of optimality, we state in passing a theorem on the existence of solutions to Problem 4.2.
@@
Suppose that there exists at least one admissible curve for Problem 4.2. Then, there exists an absolute minimizer $(K_{t},\Phi_{t})_{t\in[0,1]}$ for Problem 4.2.
Similar to Theorem 2, the proof of the above statement consists of verifying that the conditions of [9, Theorem 11.4.v.i] hold. We now proceed to state the necessary conditions of optimality for Problem 4.2.
@@
The minimizer $(\Phi_{t},K_{t})$ of Problem 4.2, or equivalently Problem 4.2, is such that
$\dot{\Phi}_{t}$ $=K_{t}\Phi_{t},$ (21a)
with $\Phi_{0}=I$ and $K_{t}$ is given by (17) and
$R_{t}$ $=-(M_{t}B+P_{t}\Phi_{t}^{\top})N_{t}^{\top}(N_{t}BN_{t}^{\top})^{-1},$ (21b)
where $\text{where }\hskip 10.0pt\dot{P_{t}}$ $=-{K_{t}}^{\top}(K_{t}\Phi_{t}\Sigma_{0}+P_{t}),$ (21c)
and $\text{and }\hskip 10.0ptB$ $=\big{(}\Phi_{t}\Sigma_{0}\Phi_{t}^{\top}+\varepsilon I\big{)}.$ (21d)
@@
then (21a) and (21c) are subject to the mixed boundary conditions $\Phi_{1}\in\mathcal{S}$ and ${P_{1}}\perp T_{\Phi_{1}}\mathcal{S}$, respectively.
Proof 4.4.
As before, we express the control Hamiltonian of Problem 4.2 as
$\mathcal{H}$ $=\frac{1}{2}{\rm tr}\big{(}K_{t}\Phi_{t}\Sigma_{0}\Phi^{\top}_{t}K_{t}^{\top}+\varepsilon K_{t}K_{t}^{\top}+P_{t}^{\top}\big{(}K_{t}\Phi_{t}\big{)}\big{)},$
for $K_{t}$ satisfying (17) and for $P_{t}$ an $n\times n$ Lagrange multiplier matrix enforcing (3). Equations (21a) and (21c) are obtained respectively from the necessary conditions $\dot{\Phi}_{t}=\partial\mathcal{H}/\partial P_{t}$ and $\dot{P}_{t}=-\partial\mathcal{H}/\partial\Phi_{t}$. Exploiting the form in (17), we rewrite the previous Hamiltonian
$\mathcal{H}$ $=\frac{1}{2}{\rm tr}\big{(}(M_{t}+R_{t}N_{t})B(M_{t}+R_{t}N_{t})^{\top}\big{)}$
@@
for $B$ satisfying (21d). Indeed, direct application of Pontryagin’s maximum principle, $\partial H/\partial R_{t}=0$, results in (21b). Lastly, Equations (10) and (5) define the surface in (21e) and conditions on $\Phi_{1}$ and $P_{1}$ from the proof of Proposition 3 apply verbatim here, which concludes the proof.
@@ 5 Examples
Figure 1: The gray envelope represents the given $\Sigma_{t\in[0,1]}$ as in (23). The green curve is the resulting trajectory of the tracer between two given endpoints (22).
@@
for $I$ denoting the identity matrix and $t\in[0,1]$. The flow $\Sigma_{t}$ corresponds to the optimal mass transport between $\Sigma_{0}$ and $\Sigma_{1}$ (McCann displacement interpolation [1]) with no data on tracer particles. This is depicted in Fig. 3, where the gray cylinder represents the evolution of a equi-probability level set of the corresponding Gaussian distributions. The availability of data on the terminal disposition of a tracer particle reveals a rotation component of the flow that cannot be gleaned from the McCann flow in (23b).
Indeed, since $\Phi_{1}{\Phi_{1}}^{\top}=\Sigma_{1}$, $\Phi_{1}=\Sigma_{1}^{\frac{1}{2}}U$, for some $U\in\rm{SO}(2)$, the special orthogonal group of dimension $2$. In the present case where $n=2$, the matrix $U$ is fully determined by the information provided. Therefore, one can solve for $U$ satisfying $\Sigma_{1}^{-\frac{1}{2}}Y_{1}=UY_{0}$, for $Y_{0},Y_{1}$ as in (22). This results in
$U=\begin{bmatrix}\frac{1}{2}\sqrt{2-\sqrt{2}}&\frac{1}{2}\sqrt{\sqrt{2}+2}\\ -\frac{1}{2}\sqrt{\sqrt{2}+2}&\frac{1}{2}\sqrt{2-\sqrt{2}}\end{bmatrix},$
@@
$\Phi_{1}=\begin{bmatrix}0&\sqrt{2}\\ -1&1\end{bmatrix},$ (24)
and the system (13a)-(13b) can be solved as a two-point boundary value problem between $\Phi_{0}=I$ and (24). The numerically obtained state-transitions $\Phi_{t}$ from the shooting method are shown in Fig. 3. These state transitions result in the trajectory depicted in green in Fig. 3, indeed reconciling the information in (22) and (23).
Example II
In contrast to the previous example, we here assume the full trajectory of a single tracer is given as shown in Fig. 6 in green. This is precisely a trajectory between the endpoints
$Y_{0}=\begin{bmatrix}-\frac{\sqrt{3}}{2}\\ \frac{1}{\sqrt{2}}\end{bmatrix},\ Y_{1}=\begin{bmatrix}0\\ -\sqrt{3}\end{bmatrix}.$ (25)
@@
The task here is to numerically demonstrate the solution of Problem 4.2 in the two-dimensional case. Similar to the previous example, the terminal state transition matrix can be fully determined as $\Phi_{1}=\Sigma_{1}^{\frac{1}{2}}O$, for
$O=\begin{bmatrix}-\frac{1}{2}&\frac{\sqrt{3}}{2}\\ \frac{\sqrt{3}}{2}&-\frac{1}{2}\end{bmatrix}\in\rm{SO}(2),$
which amounts to a rotation of $-120^{\circ}$. The resulting $\Phi_{1}$ along with $\Phi_{0}$ are used to solve the two-point boundary problem associated with system (21a)-(21d) for $\varepsilon=1$. The obtained entries of $\Phi_{t}$ are presented in Fig. 6. The gray envelope shown in Fig. 6 delineates the equi-probability level sets of the corresponding Gaussian distributions.

### v2

@@ 3.2 Internal Dynamics from Tracer Trajectories
This subsection deals with a counterpart of Problem 1. In it, the data of the problem we consider here is in the form of the distributional constraints (4) with $\mathcal{T}_{1}=\{0,1\}$ and the full trajectories of the tracers (5) with $\mathcal{T}_{2}=[0,1]$. Our task is once again to identify the dynamics in (3). If (7) is taken as the cost of transport, then a candidate problem formulation can be stated as follows.
Problem $3.$
Given $(\Sigma_{t})_{t\in\mathcal{T}_{1}}\in{\rm Sym}^{+}(n)$ with $\mathcal{T}_{1}=\{0,1\}$ and $(Y_{t})_{t\in\mathcal{T}_{2}}\in\mathbb{R}^{n\times m}$ with $\mathcal{T}_{2}=[0,1]$, find
$(K_{t}^{\star},\Phi_{t}^{\star}):=\arg\min_{(K_{t},\Phi_{t})}J_{\rm KE}$
subject to the constraints (3), (4), and (5).
@@
However, this candidate formulation may not admit a solution, in general; there are cases where $\Phi_{t}$ becomes singular at some point in time, see [22, Example 1]. To avert this issue, we modify the cost in Problem 3.3 by augmenting it with the “weighted attention” term in (9), which penalizes the $L_{2}$-norm of the gain $K_{t}$. Such an augmentation traces back to R.W. Brockett [13], according to whom,
“… the easiest control law to implement is a constant input. Anything else requires some attention. The more frequently the control changes, the more effort it takes to implement it. […], it can be argued that the cost of implementation is linked to the rate at which the control changes with changing values of $x$ [state] and $t$ [time].”
@@
For convenience, we restrict ourselves to a spatial attention term as in (8) or (9). Thereby, we state the adaptation of Problem 3.3 as follows.
Problem 3.4.
Given $(\Sigma_{t})_{t\in\mathcal{T}_{1}}\in{\rm Sym}^{+}(n)$ with $\mathcal{T}_{1}=\{0,1\}$ and $(Y_{t})_{t\in\mathcal{T}_{2}}\in\mathbb{R}^{n\times m}$ with $\mathcal{T}_{2}=[0,1]$, find
$(K_{t}^{\star},\Phi_{t}^{\star}):=\arg\min_{(K_{t},\Phi_{t})}J_{\rm KE}+\varepsilon J_{\rm A},$ (14)
@@
A curve $(K_{t},\Phi_{t})_{t\in[0,1]}$ is said to be admissible for Problem 3.4 if $(K_{t})_{t\in[0,1]}$ is square integrable and $(\Phi_{t})_{t\in[0,1]}$ is absolutely continuous, and are such that (3) is satisfied (whenever $\dot{\Phi}_{t}$ exists) as well as constraints (4) and (5), and the value of the functional (14) is finite. The following proposition provides a more explicit characterization of the class of admissible curves.
@@
Then, a curve $(K_{t},\Phi_{t})_{t\in[0,1]}$ is an admissible curve for Problem 3.4 if and only if the following four conditions hold. First, the constraint (4) is satisfied, second,
@@
third, there exists an almost everywhere unique square integrable curve $(R_{t})_{t\in[0,1]}\in\mathbb{R}^{n\times(n-m)}$ such that
@@
Proof 3.5.
Let $(K_{t},\Phi_{t})_{t\in[0,1]}$ be an admissible curve for Problem 3.4. Since $\mathcal{T}_{2}=[0,1]$ and $(Y_{t})_{t\in\mathcal{T}_{2}}$ is continuously differentiable by assumption, direct differentiation of (5) followed by substitution of (3) shows that $(K_{t})_{t\in[0,1]}$ must satisfy
@@
where $(Z_{t})_{t\in[0,1]}\in\mathbb{R}^{n\times n}$ is any square integrable curve satisfying $Z_{t}Y_{t}=0$, i.e., the column space of $Y_{t}$ is a subspace of the null space of $Z_{t}$.
To proceed, let $N_{0}\in\mathbb{R}^{(n-m)\times n}$ be any full row-rank matrix such that $N_{0}Y_{0}=0$ and define the curve $(N_{t})_{t\in[0,1]}$ as the solution to
@@
From the above, it follows that any square integrable curve $(Z_{t})_{t\in[0,1]}\in\mathbb{R}^{n\times n}$ satisfying the constraint $Z_{t}Y_{t}=0$ must be of the form $Z_{t}=R_{t}N_{t}$, for an almost everywhere unique square integrable curve $(R_{t})_{t\in[0,1]}\in\mathbb{R}^{n\times(n-m)}$. Finally, since $1\in\mathcal{T}_{2}=[0,1]$, the curve $(\Phi_{t})_{t\in[0,1]}$ must be such that $\Phi_{1}$ satisfies (16). The converse implication of the proposition follows by direct integration.
In essence, Equation (15) introduces a matrix $N_{0}$ whose null space contains the column space of $Y_{0}$. If the dynamics of $N_{t}$ coincide with (17c), then the column space of $Y_{t}$ is ensured to remain in the null space of $N_{t}$ for all $t\in[0,1]$. As such, Proposition 4 suggests a reformulation of Problem 3.4, replacing condition (5) on the state transition $\Phi_{t}$ with an explicit constraint on the gain $K_{t}$. The advantage of the new formulation lies in the fact that the search for the optimal curve reduces to searching over a lower-dimensional control space, namely, the curve $(R_{t})_{t\in[0,1]}$.
Problem 3.4
Given $(\Sigma_{t})_{t\in\mathcal{T}_{1}}\in{\rm Sym}^{+}(n)$ with $\mathcal{T}_{1}=\{0,1\}$ and $(Y_{t})_{t\in\mathcal{T}_{2}}\in\mathbb{R}^{n\times m}$ with $\mathcal{T}_{2}=[0,1]$, find
$(K_{t}^{\star},\Phi_{t}^{\star}):=\arg\min_{(K_{t},\Phi_{t})}J_{\rm KE}+\varepsilon J_{\rm A},$
for $\varepsilon>0$ and subject to the constraints (3), (4), (16), and (17).
Before we proceed with deriving the necessary conditions of optimality, we state in passing a theorem on the existence of solutions to Problem 3.4.
@@
Suppose that there exists at least one admissible curve for Problem 3.4. Then, there exists an absolute minimizer $(K_{t}^{\star},\Phi_{t}^{\star})_{t\in[0,1]}$ for Problem 3.4.
Similar to Theorem 2, the proof of the above statement consists of verifying that the conditions of [23, Theorem 11.4.vi] hold. We now proceed to state the necessary conditions of optimality for Problem 3.4.
@@
The absolute minimizer $(K_{t}^{\star},\Phi_{t}^{\star})_{t\in[0,1]}$ of Problem 3.4, or equivalently Problem 3.4, is such that
$\dot{\Phi}_{t}^{\star}$ $=K_{t}^{\star}\Phi_{t}^{\star},$ (21a)
with $\Phi_{0}^{\star}=I$ and $K_{t}^{\star}$ coincides with (17) for
$\hskip-20.0ptR_{t}$ $=-(M_{t}B+P_{t}^{\star}(\Phi_{t}^{\star})^{\top})N_{t}^{\top}(N_{t}BN_{t}^{\top})^{-1},$ (21b)
$\hskip-10.0pt\text{where }\hskip 5.0pt\dot{P}_{t}^{\star}$ $=-{(K_{t}^{\star})}^{\top}(K_{t}^{\star}\Phi_{t}^{\star}\Sigma_{0}+P_{t}^{\star}),$ (21c)
$\hskip-20.0pt\text{and }\hskip 10.0ptB$ $=\big{(}\Phi_{t}^{\star}\Sigma_{0}(\Phi_{t}^{\star})^{\top}+\varepsilon I\big{)}.$ (21d)
@@
then (21a) and (21c) are subject to the mixed boundary conditions $\Phi_{1}^{\star}\in\mathcal{S}$ and ${P_{1}^{\star}}\perp T_{\Phi_{1}^{\star}}\mathcal{S}$, respectively.
Proof 3.6.
As before, we express the control Hamiltonian of Problem 3.4 as
$\mathcal{H}(K_{t},\Phi_{t},P_{t})$ $=\frac{1}{2}{\rm tr}\big{(}K_{t}\Phi_{t}\Sigma_{0}\Phi^{\top}_{t}K_{t}^{\top}+\varepsilon K_{t}K_{t}^{\top})$
$+{\rm tr}(P_{t}^{\top}K_{t}\Phi_{t}\big{)},$
for $K_{t}$ satisfying (17) and for $P_{t}$ an $n\times n$ Lagrange multiplier matrix that enforces (3). Equations (21a) and (21c) are obtained respectively from the necessary conditions $\dot{\Phi}_{t}^{\star}=\partial\mathcal{H}/\partial P_{t}$ and $\dot{P}_{t}^{\star}=-\partial\mathcal{H}/\partial\Phi_{t}$. Exploiting the form in (17), we rewrite the previous Hamiltonian as
$\mathcal{H}(R_{t},\Phi_{t},P_{t})$ $=\frac{1}{2}{\rm tr}\big{(}(M_{t}+R_{t}N_{t})B(M_{t}+R_{t}N_{t})^{\top}\big{)}$
@@
for $B$ satisfying (21d). Indeed, direct application of Pontryagin’s maximum principle, $\partial\mathcal{H}/\partial R_{t}=0$, results in (21b). Lastly, Equations (10) and (5) define the surface in (21e) and conditions on $\Phi_{1}^{\star}$ and $P_{1}^{\star}$ from the proof of Proposition 3 apply verbatim here, which concludes the proof.
@@ 4.1 Example I
Figure 1: The gray envelope represents the given $(\Sigma_{t})_{t\in[0,1]}$ as in (23). The green curve is the resulting trajectory of the tracer between two given endpoints (22).
@@
for $t\in[0,1]$. The flow $\Sigma_{t}$ corresponds to the optimal mass transport between $\Sigma_{0}$ and $\Sigma_{1}$ (McCann’s interpolation [14, 8]) with no data on tracer particles. This is depicted in Fig. 3, where the gray cylinder represents the evolution of an equi-probability level set of the corresponding Gaussian distributions. The availability of data on the terminal disposition of a tracer particle reveals a rotational component of the flow that cannot be gleaned from the McCann flow in (23b).
Since $\Phi_{1}^{\star}{(\Phi_{1}^{\star})}^{\top}=\Sigma_{1}$, $\Phi_{1}^{\star}=\Sigma_{1}^{\frac{1}{2}}U_{1}$, for some $U_{1}\in\rm{SO}(2)$, the special orthogonal group of dimension $2$. In the present case, where $n=2$, the matrix $U_{1}$ is fully determined by the information provided. Therefore, one can solve for $U_{1}$ satisfying $\Sigma_{1}^{-\frac{1}{2}}Y_{1}=U_{1}Y_{0}$, for $Y_{0},Y_{1}$ as in (22). This results in
$U_{1}=\begin{bmatrix}\frac{1}{2}\sqrt{2-\sqrt{2}}&\frac{1}{2}\sqrt{\sqrt{2}+2}\\ -\frac{1}{2}\sqrt{\sqrt{2}+2}&\frac{1}{2}\sqrt{2-\sqrt{2}}\end{bmatrix}$,
@@
$\Phi_{1}^{\star}=\begin{bmatrix}0&\sqrt{2}\\ -1&1\end{bmatrix}$, (24)
and the system (13a)-(13c) can be solved as a two-point boundary value problem between $\Phi_{0}^{\star}=I$ and (24). The transitions $\Phi_{t}^{\star}$ shown in Fig. 3 are obtained from the shooting method.  [^1] The source code is available at:
https://github.com/a-eld/TransportWithTracers-1.
These state transitions result in the trajectory depicted in green in Fig. 3, indeed reconciling the information in (22) and (23).
4.2 Example II
In contrast to the previous example, we assume here that the full trajectory of a single tracer is given, as shown in Fig. 6 in green. This is precisely a trajectory between the endpoints:
$Y_{0}=\begin{bmatrix}-\frac{\sqrt{3}}{2}\\ \frac{1}{2}\\[3.0pt] \end{bmatrix},\ Y_{1}=\begin{bmatrix}0\\ -\sqrt{3}\end{bmatrix}$. (25)
@@ 4.2 Example II
The task here is to numerically demonstrate the solution of Problem 3.4 in the two-dimensional case. Similar to the previous example, the terminal state transition matrix can be fully determined as $\Phi_{1}^{\star}=\Sigma_{1}^{\frac{1}{2}}O_{1}$, for
$O_{1}=\begin{bmatrix}-\frac{1}{2}&-\frac{\sqrt{3}}{2}\\[3.0pt] \frac{\sqrt{3}}{2}&-\frac{1}{2}\end{bmatrix}\in\rm{SO}(2),$
which amounts to a rotation of $120^{\circ}$. The resulting $\Phi_{1}^{\star}$ along with $\Phi_{0}^{\star}=I$ are used to solve the two-point boundary value problem associated with system (21a)-(21d) for $\varepsilon=1$ and
$N_{0}=\begin{bmatrix}\frac{1}{2}&\frac{\sqrt{3}}{2}\end{bmatrix}$.
The obtained entries of $\Phi_{t}^{\star}$ are presented in Fig. 6.222The state transitions are obtained using MATLAB’s solver bvp5c. The associated source code can be found at:
https://github.com/a-eld/TransportWithTracers-2.
The gray envelope shown in Fig. 6 delineates the equi-probability level sets of the corresponding Gaussian distributions.